OpenStudy (anonymous):

pls find the equation of the parabola in standard form and identify the vertex, focus and directrix :) y^2 + 25 = 6x -1

OpenStudy (anonymous):

y^2=6x-26 y^2=6(x-(26/6)) y^2=6(x-(13/3)) \[(y-0)^{2}=4(6/4)(x-13/3)\]

OpenStudy (anonymous):

what if it there was a y already ? like it's gonna be y^2 + 20y = 6x -1

OpenStudy (anonymous):

y^2+20y+100-100=6x-1 (y+10)^2=6x-1+100 (y+10)^2=6x+99 (y+10)^2=6(x+(99/6)) (y+10)^2=6(x+33/2)

OpenStudy (anonymous):

or for standard form:::: (y+10)^2=4(6/4)(x+33/2)

OpenStudy (anonymous):

i used completing the square method out here.......

OpenStudy (anonymous):

thank you so much ! :)